The perimeter of a rectangular playground area is 308 feet. If the length of the playground is 34 feet more than twice the width, find the length and width of the playground. Construct an algebraic equation with a variable and solve using algebra steps.
Answer by checkley77(12844)

L=2W+34
2L=2W=308
2(2W+34)+2W=308
4W+68+2W=308
6W=308-68
6W=240
W=240/6
W=40 ANS. FOR THE WIDTH.
L=2*40+34
L=80+34
L=114 ANS. FOR THE LENGTH.
PROOF:
2*114+2*40=308
228+80=308
308=308